precal application problem ***please help me!***

One freezer costs $623.95 and uses a 90 kilowatt hours (kwh)of electricity each month. A second freezer costs $500 and uses 100kwh of electricity each month. Theexpected life of each freezer is12 years. What is the minimum electric rate in cents per kwh forwhich the 12-year total cost (purchase price + electricitycosts)will be less for the first freezer?***Has something to do with inequalities, but I'm just reallystuck.
Answer #1

First you need to assume the freezers are never unplugged, otherwise you will get a different answer.

Over 12 years, you have 144 months.

Assume electric rate = Y

Cost of first freezer = initial cost plus monthly cost = $723.95 + 90 kwh * 144 * Y

Cost of second freezer = initial cost plus monthly cost = $600 + 100 kwh * 144 * Y

These two costs are equal when 723.95 + 12960Y = 600 + 14400Y or 123.95 = 1440Y or Y = 8.61 cents, therefore this is the minimum cost where freezer 1 is less.